OpenStudy (anonymous):

2x – 7 = 13 7 – 3x = 9 + x 1/5x = 2x + 3

OpenStudy (solomonzelman):

2x – 7 = 13 add 7 to both sides and then divide both sides by 2 7 – 3x = 9 + x subtract 7 from both sides, subtract x from both sides, divide both sides by 2. 1/5x = 2x + 3 subtract 2x from both sides, both sides times -1 4/5 (same as times -9/5)